OpenStudy (anonymous):

Give all solutions to the following non-linear system of equations. y = x 2 − 4 y = − x 2 + 5

OpenStudy (anonymous):

is that 2x or x^2?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

so sorry about that reply

OpenStudy (anonymous):

I thought it was a good reply why did you delete it?

OpenStudy (anonymous):

y? y it gone

OpenStudy (anonymous):

sorry there are only 2 answers and I screwed up

OpenStudy (anonymous):

try it again

OpenStudy (anonymous):

sorry about that so by addition you get \[2y=1\] thus \[y=1/2\] There are two answers since both equations have a high power of 2 so y=1/2 and y=1/2 now both equations are \[x=9/2\] so \[x=3\sqrt{2}/2\] and \[x=-3\sqrt{2}/2\] so the answers are (\[3\sqrt{2}/2\],1) and (\[-3\sqrt{2}/2\],1) there's your answer, sorry the formatting is screwed up

OpenStudy (anonymous):

Thank bro!

OpenStudy (anonymous):

np bro